OpenStudy (anonymous):

How to simplify this?? \[ \frac{-\sqrt{3} - 1}{1+(-\sqrt{3})*1}\] \[ \frac{-\sqrt{3} - 1}{1-\sqrt{3}}\] \[ \frac{-\sqrt{3} - 1}{1-\sqrt{3}} * \frac{1-\sqrt{3}} {1-\sqrt{3}} \] \[ \frac{2}{2(2-\sqrt{3})} \] Is this correct??

OpenStudy (anonymous):

In step 3, you should multiply the expression by the denominator's conjugate, top and bottom.

OpenStudy (anonymous):

That allows you to eliminate radical expressions in the denominator.

OpenStudy (anonymous):

Could you show me? I don't see what you mean. I thought I did multiply the expression

OpenStudy (anonymous):

\[\frac{ -\sqrt{3}-1 }{ 1-\sqrt{3} } \times \frac{ 1+\sqrt{3} }{ 1+\sqrt{3} } \]

OpenStudy (anonymous):

That should be your step 3.

OpenStudy (anonymous):

I did that and I get 2 for the numerator and \( 2(2-\sqrt{3})\) for the denominator

OpenStudy (anonymous):

No, if you multiply a radical expression by its conjugate, the square roots disappear.

OpenStudy (anonymous):

Yeah, I used the same I see that.

OpenStudy (anonymous):

I use - and not the conjugate

OpenStudy (anonymous):

\[(1-\sqrt{3})(1+\sqrt{3}) = 1-\sqrt{3} + \sqrt{3} - 3 = 2\]

OpenStudy (anonymous):

Do you agree with my above expansion?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok, so 2 is now your denominator.

OpenStudy (anonymous):

Let me do the top

OpenStudy (anonymous):

Let's multiply the numerator now.

OpenStudy (anonymous):

I got -2

OpenStudy (anonymous):

One sec I made a boo boo

OpenStudy (anonymous):

Check your numerator again. It should contain some square roots.

OpenStudy (anonymous):

I have \[ -\sqrt{3} -\sqrt{9} -1 -\sqrt{3} \] is this correct?

OpenStudy (anonymous):

Yes, you can combine the terms to make it easier to read.

OpenStudy (anonymous):

\[-2\sqrt{3}-4\]

OpenStudy (anonymous):

Now that you have the numerator, don't forget that you still have a denominator, -2.

OpenStudy (anonymous):

From what I put up where di the 4 come from?

OpenStudy (anonymous):

Oh your \[- \sqrt{9} - 1 = -3 - 1 = - 4\]

OpenStudy (anonymous):

Combine your numerator and your denominator, now what do you have?

OpenStudy (anonymous):

Ok how about eh 2 and the sqr(3) ?? from what I put up.

OpenStudy (anonymous):

Oh that's from \[-\sqrt{3}-\sqrt{3} = -2\sqrt{3}\]

OpenStudy (anonymous):

\[\frac{ -2\sqrt{3}-4 }{ -2 } = \sqrt{3} + 2\]

OpenStudy (anonymous):

Yep that is it. Thank you so much.

OpenStudy (anonymous):

Are you a qualified helper?

OpenStudy (anonymous):

I just like to help :)